. Jones and Brown start from two points, which are 375 miles apart and
travel toward each other. The latter travels twice as fast as the former.
They meet in 5 hours. Find the rates per hour.

Answer by richwmiller(17219)

5x+5*2x=375
5*3x=375
15x=375
x=25
2x=50
5*25=125
5*50=250